A nut mixture contains 10% cashews. How many kilograms of the mixture and how many kilograms of cashews should be mixed to give 6 Kg of a nut mixture which is 25% cashews?
Answer by ankor@dixie-net.com(22740)

A nut mixture contains 10% cashews. How many kilograms of the mixture and how many kilograms of cashews should be mixed to give 6 Kg of a nut mixture which is 25% cashews?
:
Let x = the amt of 100% cashews; then (6-x) = amt of 10% mixture
:
Write a per cent cashews equation:
1.0x + .10(6-x) = .25(6)
:
1x + .6 - .1x = 1.5
:
1x - .x = 1.5 - .6
:
.9x = .9
:
x = .9/.9
:
x = 1 kg of the 100% cashews required
:
6 - 1 = 5 kg of the 10% mixture
:
Check:
1(1) + .1(5) = .25(6)
1 + .5 = 1.5